Jane’s gross monthly income is $3,300. She is applying for a mortgage loan that will have a $800 monthly mortgage payment (principal and interest). In addition, property taxes will be $200 per month, and homeowner’s insurance will be $50 per month. Jane’s other debt payments are $150 (car loan) and $100 (minimum payments on credit cards.) Will she qualify for a conventional mortgage loan if her lender requires the mortgage debt service ratio to be no more than 30% and the debt payment ratio to be no more than 40%? Hint: See Demonstration Problem 6.9.

Group of answer choices

Yes

No, her debt payment ratio is too high

No, her mortgage debt service ratio is too high

No, both ratios are too high
Answer #1

Answer :

Mortgage debt service ratio = Mortgage debt payment / Gross monthly income = 800 / 3300 = 24.24%
Debt payment ratio = Debt payment / Gross monthly income = 800 + 150 + 100 / 3300 = 31.82%

Hence 1st option "Yes" is correct as her both ratio are with limit

Kindly Up -vote Thank You !!!!